### v4

@@ Abstract
The growing Synthetic Aperture Radar (SAR) data has the potential to build a foundation model through Self-Supervised Learning (SSL) methods, which can achieve various SAR Automatic Target Recognition (ATR) tasks with pre-training in large-scale unlabeled data and fine-tuning in small labeled samples. SSL aims to construct supervision signals directly from the data, which minimizes the need for expensive expert annotation and maximizes the use of the expanding data pool for a foundational model. This study investigates an effective SSL method for SAR ATR, which can pave the way for a foundation model in SAR ATR. The primary obstacles faced in SSL for SAR ATR are the small targets in remote sensing and speckle noise in SAR images, corresponding to the SSL approach and signals. To overcome these challenges, we present a novel Joint-Embedding Predictive Architecture for SAR ATR (SAR-JEPA), which leverages local masked patches to predict the multi-scale SAR gradient representations of unseen context. The key aspect of SAR-JEPA is integrating SAR domain features to ensure high-quality self-supervised signals as target features. Besides, we employ local masks and multi-scale features to accommodate the various small targets in remote sensing. By fine-tuning and evaluating our framework on three target recognition datasets (vehicle, ship, and aircraft) with four other datasets as pre-training, we demonstrate its outperformance over other SSL methods and its effectiveness with increasing SAR data. This study showcases the potential of SSL for SAR target recognition across diverse targets, scenes, and sensors.
@@ 1 Introduction
Synthetic aperture radar (SAR) possesses the capability to capture images under diverse weather and lighting conditions, rendering it indispensable for acquiring Earth observation information (Sun et al., 2021; Moreira et al., 2013; Tsokas et al., 2022). As a crucial component in SAR image interpretation, SAR Automatic Target Recognition (ATR) aims to localize and classify desired targets within SAR images. Over the past few decades, extensive research has been conducted in this field with various civilian and military applications, including transportation management (Gagliardi et al., 2023), autonomous driving (Rizzi et al., 2021), and military surveillance (Li et al., 2023b; Peng et al., 2022). Deep learning has revitalized SAR ATR in the last decade with remarkable advancements (Kechagias-Stamatis and Aouf, 2021; Li et al., 2023a; Datcu et al., 2023). Despite significant progress, the high costs of data collection and annotation limit dataset diversity and method applications. SAR ATR studies are categorized into subfields such as vehicle (Li et al., 2023b), ship (Hou et al., 2020), and aircraft recognition (Zhao et al., 2022) with many specialized approaches. Hence, an intriguing and worthwhile question to explore is whether a generic representation and method exist for various SAR target recognition tasks.
@@
The emerging self-supervised learning techniques present a potential solution to this question. Self-supervised learning (SSL) aims to extract meaningful signals directly from the data, thereby reducing the need for costly expert annotations and efficiently using the growing number of data. Foundation models, which are trained on extensive datasets using SSL, have demonstrated their effectiveness in various computer vision (Goldblum et al., 2023; He et al., 2022), remote sensing (Sun et al., 2023), and clinical medical tasks (Zhou et al., 2023b). Therefore, we believe that foundation models, trained through SSL on a large dataset, can yield a generalized representation for SAR target recognition. With this aim, this paper investigates the key challenges in SSL for SAR ATR.
@@
Figure 1: Comparison of related architectures and the proposed architecture. (a) Generative architectures (He et al., 2022) focus on reconstructing the pixels of unseen patches with a high mask proportion. This approach creates a challenging and meaningful pretext task that allows the model to learn contextual relationships within images. (b) Physics-guided contrastive architectures (PGCA) (Datcu et al., 2023; Huang et al., 2022) leverage the unique representation in the SAR domain as guided signals to constrain the representation of the Deep Neural Network (DNN). By incorporating physical principles into the learning process, PGCA enhances the model’s ability to capture accurate and important SAR features. (c) Joint-Embedding Predictive Architecture (I-JEPA) (Assran et al., 2023) use deep features as target signals and learn more semantic information. However, we found that a learnable target encoder is susceptible to feature collapses due to SAR image noise. (d) Our Joint-Embedding Predictive Architecture for SAR ATR (SAR-JEPA) combines SAR domain embedding and the meaningful MIM pretext task to learn contextual relationships in the SAR gradient feature space. This approach utilizes prior knowledge about SAR target recognition in target scale and features to improve the representation.
@@
Lack of a large-scale pre-training set. Just as ImageNet (Fei-Fei and Krishna, 2022) has achieved an impressive contribution to computer vision by pushing the number of categories for object recognition from twenty to thousands, a large-scale pre-training set for SAR target recognition will bring new research ideas and tasks. Unfortunately, due to the acquisition cost and annotation difficulties, many SAR target datasets have at most ten annotation categories and a few thousand images. Whereas some studies (Huang et al., 2022; Zhai et al., 2022; Pei et al., 2023) in SAR target recognition used a few datasets as pre-training and evaluations, the potential of self-supervised learning in large-scale SAR datasets has yet to be explored fully. Therefore, it is necessary to integrate increasing open-source datasets to construct a large-scale pre-training set containing diverse targets, scenes, and sensors [^1] Various imaging conditions regarding target, scene, and sensor can lead to complex target and background signature variations in SAR images (Ross et al., 1999; Li et al., 2023b)..
@@
Special properties of SAR images. SAR achieves coherent imaging through a moving platform in remote sensing. Its images reflect the electromagnetic scattering of objects, but they are often affected by multiplicative speckle noise, phase errors, and other interferences. Consequently, SAR image interpretation is extremely difficult and poses difficulties for SSL methods due to unique imaging properties. Our study focuses on addressing two key issues: the scale problem and target features. Firstly, remote sensing images often contain large scenes and small targets, making it difficult to learn effective contextual information for various targets. Current pre-training methods for SAR target recognition focus less on this issue because they have mainly focused on a few datasets that lack significant scale problems in the dataset. Secondly, SAR images are affected by speckle noise, which can impact the quality of learned representations. This issue was addressed through data augmentation (Zhai et al., 2022) and filtering methods (Pei et al., 2023) in contrastive SSL, as well as Histogram of Oriented Gradients (HOG) features (Wang et al., 2023d) in generative SSL. However, data augmentation in contrastive learning can not simulate the effects of speckle noise exactly, and the gradient computation in HOG features has false results due to multiplicative noise. Based on the above discussions, these two issues still require adequate solutions. We aim to leverage SAR domain knowledge to address these unique challenges for SSL in pretext tasks and target signals.
@@
In order to deploy SSL for SAR ATR, we first constructed a pre-training dataset consisting of approximately 100,000 SAR magnitude images sourced from four open-source datasets (Xia et al., 2022; Chen et al., 2022b; Wang et al., 2019; Malmgren-Hansen et al., 2017; Kusk et al., 2016; Lewis et al., 2019). This pre-training dataset encompasses a wide range of target categories, scenes, and sensors, allowing us to explore effective SSL methods. To address the abovementioned issues, we proposed a novel Joint-Embedding Predictive Architecture for SAR ATR (SAR-JEPA [^2] Our codes and weights: https://github.com/waterdisappear/SAR-JEPA) in Figure 1. SAR-JEPA leverages gradient embedding and the Masked Image Modeling (MIM) task to learn contextual relationships within the SAR feature space. We employed the well-established Gradient-by-Ratio (GR) method (Dellinger et al., 2014; Song et al., 2016) as target feature, effectively mitigating the impact of speckle noise in SAR images and facilitating accurate target shape extraction. Furthermore, we introduced local masks and multi-scale features to enhance the architecture’s adaptability to various small target scales in remote sensing. To evaluate the effectiveness of our proposed architecture, we conducted experiments on three datasets (vehicle, ship, and aircraft).
@@
Numerous experiment results demonstrate that SAR-JEPA achieves improved performance than other SSL methods. Importantly, it can increase with the volume of pre-training data, thereby showing the potential of SSL to build a foundation model to unify SAR ATR across different targets, scenes, and sensors. In summary, the main contributions of this article are as follows:
@@ 1. $\bullet$
    We construct comprehensive SAR pre-training and fine-tuning settings to study SSL for SAR ATR self-supervised learning. A novel joint-embedding predictive architecture for SAR ATR (SAR-JEPA) is proposed to learn contextual relationships within the SAR gradient feature space.
@@ 2. $\bullet$
    The core of SAR-JEPA lies in the self-supervised learning method and the target feature construction. SAR-JEPA learns contextual information around small targets in remote sensing images more efficiently by local masking. SAR-JEPA uses multi-scale gradient features as guidance signals to solve SAR speckle noise interference.
@@ 3. $\bullet$
    SAR-JEPA’s performance with increasing data volume demonstrates the potential of SSL to learn generalized feature representations. We hope it will stimulate the enthusiasm for research on fundamental models of SAR target recognition.
The remainder of this paper is organized as follows. Sec. 2 introduces related studies in SSL. Sec. 3 introduces the proposed SAR-JEPA. Sec. 4 conducts extensive experiments and analyses for the effectiveness and scaling capabilities of our architecture. Sec. 5 concludes the whole paper and discusses future work.
@@ 2.1 Self-supervised learning in computer vision
Self-supervised learning (Liu et al., 2021; Balestriero et al., 2023) aims to learn the intrinsic relationship in samples through a pretext task. This method can leverage a huge amount of unlabelled data and perform well in various downstream tasks. Self-supervised learning methods can be broadly categorized into contrastive and generative approaches (Liu et al., 2021). In this study, we focus on generative masked image modeling because it can model both local and long-range information (Xie et al., 2023b), and the data augmentation pretext task in the contrastive approach is not fully compatible with the SAR image properties. BEiT (Bao et al., 2021) proposed the concept of masked image modeling and predicted the masked patches in a frozen pre-training tokenized space. Compared to BEiT, MAE (He et al., 2022) achieved an efficient and meaningful task through asymmetric autoencoder structure and high masking rate. Furthermore, SimMIM (Xie et al., 2022) showed that masked image modeling can be achieved with a simple structure, such as a linear projection layer and a random mask. Over time, a series of improvement works have emerged, and our focus is on research in masking methods and target features.
@@
Masking strategy can affect the representation quality, which is centered on ensuring pretext task difficulty and learning contextual information around target regions. MAE (He et al., 2022) and SimMIM (Xie et al., 2022) compared different sampling strategies, such as random sampling, block sampling, and grid sampling, and found that random sampling with a large mask ratio works best. I-JEPA (Assran et al., 2023) employed block sampling and used a single context block to predict several non-overlapping target blocks to exploit the rich semantics in the context block. Hard patches mining (Wang et al., 2023b) applied a teacher network to predict the reconstruction loss of different patches to better mine the target foreground region. PixMIM (Liu et al., 2023) used a simple random resized crop to preserve more of the target foreground region in the random mask for semantic concept learning. Inspired by the fact that target reconstruction in MAE mainly uses the surrounding area, Local Masked Reconstruction (LoMaR) (Chen et al., 2022a) proposed to perform masked reconstruction within a local image block instead of the whole image to improve computational efficiency. As for remote sensing images, the masking strategy must consider their large scene and small target properties. We applied the LoMaR method to learn local contextual information.
@@
Target features. Masked image modeling learns contextual information by reconstructing and predicting various feature representations, such as pixel values (He et al., 2022; Xie et al., 2022), discrete tokens (Bao et al., 2021), HOG features (Wei et al., 2022; Wang et al., 2023c), deep features (Wei et al., 2022; Assran et al., 2023), and frequency features (Xie et al., 2023a; Liu et al., 2023). MaskFeat (Wei et al., 2022) proved that the target features can be either the classical HOG feature or deep features obtained by other self-supervised methods. LocalMIM (Wang et al., 2023c) performed multiscale reconstruction of HOG features at different encoder layers to facilitate representation learning speed and semantic understanding of scales. I-JEPA (Assran et al., 2023) used a learnable target encoder to get deep semantic features. PixMIM (Liu et al., 2023) enhanced the learning of low-frequency components such as target shape by reconstructing the image after low-pass filtering. Masked frequency modeling (Xie et al., 2023a) avoided spatially redundant masking by reconstructing different frequency components. A priori knowledge can be incorporated as a target feature to guide representation learning. In particular, our SAR-JEPA is inspired by the architecture of I-JEPA, which uses a learnable target encoder to learn deep semantic information. However, we found that this approach leads to feature collapse with SAR image noise. Therefore, we used the gradient-by-ratio method in SAR to extract the target edge shape features and suppress the speckle noise.
@@ 2.2 Self-supervised learning in remote sensing
In recent years, there has been a growing interest in self-supervised learning for remote sensing, which provides a paradigm for solving the contradiction between increasing samples and the lack of high-quality annotations (Wang et al., 2022b). Self-supervised learning has many applications in multi-source data such as multispectral (Reed et al., 2023; Sun et al., 2023), hyperspectral (Ibanez et al., 2022), and SAR (Wang et al., 2023d; Huang et al., 2022; Zhai et al., 2022) with various paradigms of contrastive and generative.
@@
Due to the lack of a large-scale dataset, self-supervised learning in SAR ATR is relatively scarce (Zhai et al., 2022; Pei et al., 2023; Li et al., 2023b; Wen et al., 2021). These studies rely on contrastive learning or target rotation angle information. However, data augmentation methods applied on natural images may not be fully suitable for SAR images (Wang et al., 2022b). For example, Gaussian blurring and noise methods can not simulate multiplicative speckle noise effects, and the rotation is incompatible with SAR images because of the anisotropic electromagnetic scattering. In addition, only a few datasets provide information on the target rotation angle. Compared to applying multiple data augmentation methods to simulate the imaging condition variations and noise interference for SAR images (Zhai et al., 2022; Pei et al., 2023; Li et al., 2023b), we prefer another contrastive architecture (Datcu et al., 2023) compatible with SAR images’ properties. Its contrastive architecture used SAR domain features to guide representation learning. We combine this idea with MIM to learn contextual semantic information within the feature space. And the target features with the SAR image properties can ensure a high-quality representation. We discuss self-supervised learning in remote sensing and SAR from three aspects.
@@
Pre-train dataset. There is no unified benchmark setting for self-supervised learning with SAR target recognition due to the lack of a large dataset such as ImageNet. For scene classification (Wang et al., 2023d; Huang et al., 2022), the pre-training sets used BigEarthNet-MM (Sumbul et al., 2021) or the sea ice dataset (Huang et al., 2022). For target recognition (Zhai et al., 2022; Pei et al., 2023), the pre-training sets applied the MSTAR vehicle (Air Force Research Laboratory, ) or OpenSAR ship (Huang et al., 2017) datasets. Previous work for SAR target recognition (Zhai et al., 2022; Pei et al., 2023) used pre-training sets that contain several target categories and scenarios. Therefore, the potential of self-supervised learning for foundation models in SAR target recognition has yet to be investigated. Constructing a more large dataset containing various targets, scenes, and sensors is necessary.
@@
Scale of target and scene. A major difference between remote sensing and computer vision is in scene range and target size. Remote sensing images usually contain large scenes and small targets (Li et al., 2023c; Cheng et al., 2023). TOV (Tao et al., 2023) illustrated that small scenes with different semantics within a larger scene affect contrastive learning. Chen et al. (Chen and Bruzzone, 2022) used shift alignment of two local regions between SAR and optical images in contrast learning. RingMo (Sun et al., 2023) proposed a PIMask sampling strategy for dense small targets in remote sensing images, i.e., secondary pixel sampling of the masked patches to expose small targets. However, we found that pixel sampling does not work for SAR images because the single pixels of SAR images contain multiplicative noise. Therefore, we prefer local patches for masking instead of the whole image or pixel level.
@@
SAR speckle noise. Speckle noise affects the quality of self-supervised learning features, so the researcher has used many ways, such as data augmentation (Zhai et al., 2022), filtering (Pei et al., 2023), and feature extraction, to address the problem. Zhai et al. (Zhai et al., 2022) added Gaussian noise and blur to data augmentation methods. But Gaussian noise and blurring do not perfectly match multiplicative speckle noise. Pei et al. (Pei et al., 2023) used the despeckling pre-processing method in contrastive learning to address speckle noise. We focus on extracting high-quality features as guide signals, which accomplish noise reduction and feature extraction. Similarly to our work, Wang et al. (Wang et al., 2023d) discussed different features such as CannyEdge (Canny, 1986), HOG, and SAR-SIFT (Dellinger et al., 2014) features for SAR scene classification and segmentation. They found HOG is more appropriate than SAR-SIFT for SAR SSL. Our work aims to mine semantic information at the target level rather than the scene level through local patches, and we found that gradient-by-ratio is more suitable for target shape information extraction than HOG’s differential gradient under multiplicative speckle noise.

### v5

@@ Abstract
The growing Synthetic Aperture Radar (SAR) data can build a foundation model using self-supervised learning (SSL) methods, which can achieve various SAR automatic target recognition (ATR) tasks with pretraining in large-scale unlabeled data and fine-tuning in small-labeled samples. SSL aims to construct supervision signals directly from the data, minimizing the need for expensive expert annotation and maximizing the use of the expanding data pool for a foundational model. This study investigates an effective SSL method for SAR ATR, which can pave the way for a foundation model in SAR ATR. The primary obstacles faced in SSL for SAR ATR are small targets in remote sensing and speckle noise in SAR images, corresponding to the SSL approach and signals. To overcome these challenges, we present a novel joint-embedding predictive architecture for SAR ATR (SAR-JEPA) that leverages local masked patches to predict the multi-scale SAR gradient representations of an unseen context. The key aspect of SAR-JEPA is integrating SAR domain features to ensure high-quality self-supervised signals as target features. In addition, we employ local masks and multi-scale features to accommodate various small targets in remote sensing. By fine-tuning and evaluating our framework on three target recognition datasets (vehicle, ship, and aircraft) with four other datasets as pretraining, we demonstrate its outperformance over other SSL methods and its effectiveness as the SAR data increases. This study demonstrates the potential of SSL for the recognition of SAR targets across diverse targets, scenes, and sensors. Our codes and weights are available in https://github.com/waterdisappear/SAR-JEPA.
@@ 1 Introduction
Synthetic aperture radar (SAR) can capture images under various weather and lighting conditions, making it indispensable for acquiring Earth observation information (Sun et al., 2021; Moreira et al., 2013; Tsokas et al., 2022). As a crucial component in the interpretation of the SAR images, SAR automatic target recognition (SAR ATR) aims to localize and classify the desired targets within SAR images. Over the past few decades, extensive studies have been conducted in this field with various civilian and military applications, including transportation management (Gagliardi et al., 2023), autonomous driving (Rizzi et al., 2021), and military surveillance (Li et al., 2023b; Peng et al., 2022). Deep learning has revitalized SAR ATR in the last decade with remarkable advancements (Kechagias-Stamatis and Aouf, 2021; Li et al., 2023a; Datcu et al., 2023). Despite significant progress made, the high costs of data collection and annotation limit dataset diversity and method applications. SAR ATR studies are classified into subfields, such as vehicles (Li et al., 2023b), ships (Hou et al., 2020), and aircraft recognition (Zhao et al., 2022), using many specialized approaches. Hence, an intriguing and worthwhile question to explore is whether a generic representation and method exist for various SAR target recognition tasks.
@@
Emerging self-supervised learning (SSL) techniques present a potential answer to this question. SSL aims to extract meaningful signals directly from the data, thereby reducing the need for costly expert annotations and efficiently using an increasing amount of data. Foundation models trained on extensive datasets using SSL have demonstrated their effectiveness in various computer vision (Goldblum et al., 2023; He et al., 2022), remote sensing (Sun et al., 2023), and clinical medical tasks (Zhou et al., 2023b). Therefore, we believe that foundation models, trained using SSL on a large dataset, can produce a generalized representation for SAR target recognition. Consequently, this study investigates the key challenges in SSL for SAR ATR.
@@
Figure 1: Comparison of related architectures and the proposed architecture. (a) Generative architecture (He et al., 2022) focuses on reconstructing the pixels of unseen patches with a high mask proportion. This approach creates a challenging and meaningful pretext task that allows the model to learn contextual relationships within images. (b) Physics-guided contrastive architecture (PGCA) (Datcu et al., 2023; Huang et al., 2022) leverages the unique representation in the SAR domain as guided signals to constrain the representation of the deep neural network (DNN). By incorporating physical principles into the learning process, PGCA improves the ability of the model to capture accurate and important SAR features. (c) Image-based joint-embedding predictive architecture (I-JEPA) (Assran et al., 2023) uses deep features as target signals and learns more semantic information. However, we found that a learnable target encoder is susceptible to feature collapses due to SAR image noise. (d) Our joint-embedding predictive architecture for SAR ATR (SAR-JEPA) combines the SAR domain embedding and the meaningful masked image modeling (MIM) pretext task to learn contextual relationships in the SAR gradient feature space. This approach utilizes prior knowledge about SAR target recognition on target scale and features to improve representation.
@@
Lack of a large-scale pretraining set. Just as ImageNet (Fei-Fei and Krishna, 2022) has achieved an impressive contribution to computer vision by increasing the number of categories for object recognition from 20 to thousands, a large-scale pretraining set for SAR target recognition can introduce new research ideas and tasks. Unfortunately, owing to the acquisition cost and annotation difficulties, many SAR target datasets have at most ten annotation categories and a few thousand images. Although some studies (Huang et al., 2022; Zhai et al., 2022; Pei et al., 2023) on SAR target recognition used some datasets as pretraining and evaluations, the potential of SSL in large-scale SAR datasets has not yet been fully explored. Therefore, it is necessary to integrate increasing open-source datasets to construct a large-scale pretraining set that contains various targets, scenes, and sensors [^1] Various imaging conditions regarding target, scene, and sensor can lead to complex target and background signature variations in SAR images (Ross et al., 1999; Li et al., 2023b)..
@@
Special properties of SAR images. SAR achieves coherent imaging using a moving platform in remote sensing. Its images reflect the electromagnetic scattering of objects, but they are often affected by multiplicative speckle noise, phase errors, and other interferences. Consequently, SAR image interpretation is extremely difficult and poses difficulties for SSL methods because of their unique imaging properties. Our study focuses on addressing two key issues: the scale problem and target features. First, remote sensing images often contain large scenes and small targets, making it difficult to learn effective contextual information for various targets. Current pretraining methods for SAR target recognition focus insignificantly on this issue because they mainly use a few datasets that lack significant scale problems in the dataset. Second, SAR images are affected by speckle noise, which can affect the quality of learned representations. This issue was addressed through data augmentation (Zhai et al., 2022) and filtering methods (Pei et al., 2023) in contrastive SSL, as well as the histogram of oriented gradient (HOG) features (Wang et al., 2023d) in generative SSL. However, data augmentation in contrastive learning cannot exactly simulate the effects of speckle noise, and the gradient computation in HOG features has false results due to multiplicative noise. Based on the above discussions, these two issues still require adequate solutions. We aim to leverage the knowledge of the SAR domain to address these unique challenges of SSL in pretext tasks and target signals.
@@
To deploy SSL for SAR ATR, we first constructed a pretraining dataset comprising approximately 100,000 SAR magnitude images sourced from four open-source datasets (Xia et al., 2022; Chen et al., 2022b; Wang et al., 2019; Malmgren-Hansen et al., 2017; Kusk et al., 2016; Lewis et al., 2019). This pretraining dataset encompasses a wide range of target categories, scenes, and sensors, allowing us to explore effective SSL methods. To address the aforementioned issues, we proposed a novel joint-embedding predictive architecture for SAR ATR (SAR-JEPA) in Figure 1. SAR-JEPA leverages gradient embedding and the masked image modeling (MIM) task to learn contextual relationships within the SAR feature space. We employed the well-established gradient-by-ratio (GR) (Dellinger et al., 2014; Song et al., 2016) method as the target feature, thereby effectively mitigating the effect of speckle noise in SAR images and facilitating accurate target shape extraction. Furthermore, we introduced local masks and multi-scale features to enhance the adaptability of the architecture to various small target scales in remote sensing. To evaluate the effectiveness of the proposed architecture, we conducted experiments on three datasets (vehicle, ship, and aircraft).
@@
Numerous experimental results demonstrate that SAR-JEPA achieves improved performance than other SSL methods. Importantly, it can increase with the volume of pretraining data, thereby showing the potential of SSL to build a foundation model to unify the SAR ATR across various targets, scenes, and sensors. The main contributions of this study are as follows.
@@ 1. $\bullet$
    We construct comprehensive SAR pretraining and fine-tuning settings to study SSL for SAR ATR self-supervised learning. A novel joint-embedding predictive architecture for SAR ATR (SAR-JEPA) is proposed to learn contextual relationships within the SAR gradient feature space.
@@ 2. $\bullet$
    The core of SAR-JEPA depends on the SSL method and target feature construction. SAR-JEPA learns contextual information around small targets in remote sensing images more efficiently by local masking. SAR-JEPA uses multi-scale gradient features as guidance signals to solve SAR speckle noise interference.
@@ 3. $\bullet$
    SAR-JEPA’s performance with an increasing data volume demonstrates the potential of SSL to learn generalized feature representations. We hope it will stimulate the enthusiasm for research on fundamental models of SAR target recognition.
The rest of this paper is organized as follows. Section 2 introduces related studies on SSL. Section 3 introduces the proposed SAR-JEPA. Section 4 presents extensive experiments and analyses of the effectiveness and scaling capabilities of our architecture. Section 5 concludes the paper and discusses future work.
@@ 2.1 Self-supervised learning in computer vision
SSL (Liu et al., 2021; Balestriero et al., 2023) aims to learn the intrinsic relationship in samples through a pretext task. This method can leverage a large amount of unlabeled data and perform well in various downstream tasks. SSL methods can be broadly classified into contrastive and generative approaches (Liu et al., 2021). In this study, we focus on generative masked image modeling because it can model both local and long-range information (Xie et al., 2023b), and the data augmentation pretext task in the contrastive approach is not fully compatible with the properties of the SAR image. Bidirectional encoder representation from image transformers (BEiT) (Bao et al., 2021) proposed the concept of masked image modeling and predicted masked patches in a frozen pretraining tokenized space. Compared to BEiT, the mask autoencoder (MAE) (He et al., 2022) achieved an efficient and meaningful task through the asymmetric autoencoder structure and the high masking rate. Furthermore, simple framework for MIM (SimMIM) (Xie et al., 2022) has shown that masked image modeling can be achieved with a simple structure, such as a linear projection layer and a random mask. Over time, several improvement studies have emerged, and our focus has been on research on masking methods and target features.
@@
Masking strategy can affect the quality of the representation, which focuses on ensuring the difficulty of the pretext task and learning contextual information around the target regions. MAE (He et al., 2022) and SimMIM (Xie et al., 2022) compared different sampling strategies, such as random sampling, block sampling, and grid sampling, and found that random sampling with a large mask ratio worked the best. I-JEPA (Assran et al., 2023) employed block sampling and used a single context block to predict several nonoverlapping target blocks to exploit the rich semantics in the context block. Hard patch mining (Wang et al., 2023b) applied a teacher network to predict the reconstruction loss of different patches to better mine the target foreground region. Pixel reconstruction in MIM (PixMIM) (Liu et al., 2023) used a simple random resized crop to preserve more of the target foreground region in the random mask for learning semantic concepts. Inspired by the fact that the target reconstruction in MAE uses mainly the surrounding area, Local masked reconstruction (LoMaR) (Chen et al., 2022a) was proposed to perform masked reconstruction within a local image block instead of the entire image to improve computational efficiency. For remote sensing images, the masking strategy must consider their large scene and small target properties. We applied the LoMaR method to learn the local contextual information.
@@
Target features. Masked image modeling learns contextual information by reconstructing and predicting various feature representations, such as pixel values (He et al., 2022; Xie et al., 2022), discrete tokens (Bao et al., 2021), HOG features (Wei et al., 2022; Wang et al., 2023c), deep features (Wei et al., 2022; Assran et al., 2023), and frequency features (Xie et al., 2023a; Liu et al., 2023). MaskFeat (Wei et al., 2022) proved that the target features can be either the classical HOG features or the deep features obtained using other self-supervised methods. MIM with local multi-scale reconstruction (LocalMIM) (Wang et al., 2023c) performed multi-scale reconstruction of HOG features at different encoder layers to facilitate the representation learning speed and semantic understanding of scales. I-JEPA (Assran et al., 2023) used a learnable target encoder to obtain deep semantic features. PixMIM (Liu et al., 2023) improved the learning of low-frequency components, such as target shape, by reconstructing the image after low-pass filtering. Masked frequency modeling (Xie et al., 2023a) avoided spatially redundant masking by reconstructing various frequency components. A priori knowledge can be incorporated as a target feature to guide the representation learning. In particular, our SAR-JEPA is inspired by the architecture of I-JEPA, which uses a learnable target encoder to learn deep semantic information. However, we found that this approach leads to feature collapse with SAR image noise. Therefore, we used the GR method in SAR to extract the features of the target shape and suppress speckle noise.
@@ 2.2 Self-supervised learning in remote sensing
In recent years, there has been increasing interest in SSL for remote sensing, which provides a paradigm for solving the contradiction between the increase in samples and the lack of high-quality annotations (Wang et al., 2022b). SSL has many applications in multi-source data, such as multispectral (Reed et al., 2023; Sun et al., 2023), hyperspectral (Ibanez et al., 2022), and SAR (Wang et al., 2023d; Huang et al., 2022; Zhai et al., 2022), with various contrastive and generative paradigms.
@@
Owing to the lack of a large-scale dataset, SSL in SAR ATR is relatively scarce (Zhai et al., 2022; Pei et al., 2023; Li et al., 2023b; Wen et al., 2021). These studies rely on contrastive learning or information on the target rotation angle. However, the data augmentation methods applied to natural images may not be fully suitable for SAR images (Wang et al., 2022b). For example, Gaussian blurring and noise methods cannot simulate multiplicative speckle noise effects, and the rotation is incompatible with SAR images because of the anisotropic electromagnetic scattering. In addition, only a few datasets provide information on the target rotation angle. Compared to applying multiple data augmentation methods to simulate imaging condition variations and noise interference for SAR images (Zhai et al., 2022; Pei et al., 2023; Li et al., 2023b), we prefer another contrastive architecture (Datcu et al., 2023) compatible with the properties of the SAR image. Its contrastive architecture used SAR domain features to guide representation learning. We combine this idea with MIM to learn the contextual semantic information within the feature space. The target features with the SAR image properties can ensure a high-quality representation. We discuss SSL learning in remote sensing and SAR from three viewpoints.
@@
Pretrain dataset. There is no unified benchmark setting for SSL with SAR target recognition owing to the lack of a large dataset, such as ImageNet. For scene classification (Wang et al., 2023d; Huang et al., 2022), the pretraining sets used BigEarthNet-MM (Sumbul et al., 2021) or the sea ice dataset (Huang et al., 2022). For target recognition (Zhai et al., 2022; Pei et al., 2023), the pretraining sets applied the moving and stationary target acquisition and recognition (MSTAR) vehicle (Air Force Research Laboratory, ) or OpenSARship (Huang et al., 2017) datasets. Previous studies on SAR target recognition (Zhai et al., 2022; Pei et al., 2023) used pretraining sets that contained several target categories and scenarios. Therefore, the potential of SSL for foundation models in SAR target recognition has not yet been investigated. Constructing a larger dataset containing various targets, scenes, and sensors is necessary.
@@
Scale of target and scene. A major difference between remote sensing and computer vision is in scene range and target size. Remote sensing images usually contain large scenes and small targets (Li et al., 2023c; Cheng et al., 2023). The original vision (TOV) (Tao et al., 2023) illustrated that small scenes with different semantics within a larger scene affect contrastive learning. SAR-optical data fusion (Chen and Bruzzone, 2022) used the shift alignment of two local regions between SAR and optical images in contrast learning. Remote sensing foundation model framework (RingMo) (Sun et al., 2023) proposed a patch incomplete mask sampling strategy for small dense targets in remote sensing images, that is, secondary pixel sampling of the masked patches to expose small targets. However, we found that pixel sampling does not work for SAR images because the single pixels in the SAR images contain multiplicative noise. Therefore, we prefer local patches for masking instead of the entire image or the pixel level.
@@
SAR speckle noise. Speckle noise affects the quality of the SSL features. To address this problem, researchers have applied many methods, such as data augmentation (Zhai et al., 2022), filtering (Pei et al., 2023), and feature extraction. Batch instance discrimination and feature clustering (BIDFC) (Zhai et al., 2022) added Gaussian noise and blur to data augmentation methods. However, Gaussian noise and blurring do not perfectly match the multiplicative speckle noise. The despeckling preprocessing was used in contrastive learning to address speckle noise (Pei et al., 2023). We focus on extracting high-quality features, such as guide signals, which accomplish noise reduction and feature extraction. Similar to our study, feature guided masked autoencoder (FG-MAE) (Wang et al., 2023d) discussed different features such as CannyEdge (Canny, 1986), HOG, and SAR scale-invariant feature transform (SAR-SIFT) (Dellinger et al., 2014) features for SAR scene classification and segmentation. They found that HOG is more appropriate than SAR-SIFT for SAR SSL. This study aims to extract semantic information at the target level rather than at the scene level through local patches, and we found that GR is more suitable for target shape information extraction than the differential gradient of HOG under multiplicative speckle noise.